# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

### v3

@@ 5.1 Leveraging the Q-function for dimension reduction
Suppose that D1 holds. Then,
$\Psi_{q_{n}}(P_{0})-\Psi(P_{0})$ $=\langle\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})-\mathcal{T}_{0,q_{n}}(\widetilde{\alpha}_{0,q_{n}}),\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}}-q_{0})\rangle_{P_{0}}.$
If D2 also holds, then $\Psi_{q_{n}}(P_{0})-\Psi(P_{0})=O_{p}\left(\|q_{n}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\|_{P_{0}}\right).$
Condition D1 is a variant of C1 for the dimension-reduced Bellman operator. Theorem 5 states that the parameter approximation bias vanishes if the optimal transformation of the estimated $Q$-function $q_{n}$ converges to $q_{0}$, and the difference between $\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})$—a function of $(q_{n}(A_{0},S_{0}),q_{0}(A_{0},S_{0}))$—and its projection onto functions of $q_{n}(A_{0},S_{0})$ also vanishes in $\|\cdot\|_{P_{0}}$. Heuristically, this requires that conditioning on $(q_{n}(A_{0},S_{0}),q_{0}(A_{0},S_{0}))$ asymptotically provides the same information as conditioning on either component alone. We formalize this by imposing D2, a mild regularity requirement that a certain bivariate function is Lipschitz continuous. Related smoothness conditions for debiased machine learning with estimated features in regression have been studied in Benkeser et al. (2017, 2020); Wang et al. (2023b); Bonvini et al. (2024); van der Laan et al. (2024b).
5.2 Proposed calibrated DRL estimator
In this section, we propose a DRL estimator of the oracle parameter $\Psi_{q_{0}}(P_{0})$, which can be interpreted as an adaptive estimator of $\Psi(P_{0})$ in the sense of Section 3.4. Instead of using an influence-function-based bias correction, we introduce a novel debiased plug-in estimator inspired by calibrated DML (van der Laan et al., 2024b). Our approach leverages a new form of calibration—Bellman calibration—which automatically corrects bias for any continuous linear functional, without requiring explicit estimation of the Riesz representer.
Our proposed Bellman-calibrated estimator of $\Psi(P_{0})$ is given by the plug-in estimator
$\psi_{n}^{*}:=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*}),$
where $q_{n}^{*}$ is a Bellman-calibrated estimator of the $Q$-function $q_{0}$ constructed to satisfy the empirical Bellman equation:
$q_{n}^{*}(a,s)=\frac{\sum_{i=1}^{n}1(q_{n}^{*}(A_{0,i},S_{0,i})=q_{n}^{*}(a,s))\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})\}}{\sum_{i=1}^{n}1(q_{n}^{*}(A_{0,i},S_{0,i})=q_{n}^{*}(a,s))}.$ (10)
As one approach to constructing a Bellman-calibrated estimator, we propose isotonic Bellman calibration, which is outlined in Algorithm 2 and discussed in more detail below. Bellman calibration implies that the Bellman residuals $\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\}_{i=1}^{n}$ are empirically orthogonal to the transformed $Q$-function estimates $\{f(q_{n}^{*}(A_{0,i},S_{0,i}))\}_{i=1}^{n}$ for each $f:\mathbb{R}\rightarrow\mathbb{R}$. A consequence of this property is the following lemma.
Lemma 6 (Calibration corrects plug-in bias).
Suppose $q_{n}^{*}$ is Bellman calibrated (10). Then,
A consequence of Lemma 6 is that the Bellman-calibrated estimator $\psi_{n}^{*}$ is implicitly debiased for the data-dependent parameter $\Psi_{q_{n}^{*}}$. In particular, it equals a DRL estimator of $\Psi_{q_{n}^{*}}$ that uses $q_{n}^{*}$ as an estimate of $q_{0}$ and the true Riesz representer nuisance $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$. By Theorem 5, it is also debiased for the oracle parameter $\Psi_{q_{0}}$, up to second-order terms. Consequently, $\psi_{n}^{*}$ is an example of an adaptive DRL estimator, as described in Section 3.4, corresponding to the working model $H_{n}=\{f\circ q_{n}^{*}:f\}$ and the oracle model $H_{0}=\{f\circ q_{0}:f\}$. The debiasing achieved via calibration is agnostic to the specific functional and does not require knowledge of $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$.
Algorithm 2 Isotonic Bellman Calibration
0:  Initial estimator $q_{n}$ of $q_{0}$, stopping threshold $\varepsilon\approx 0$;
1:  initialize $q_{n}^{*(0)}:=q_{n}$;
2:  for $k=0,1,2,\ldots$ do
3:     compute $f_{n}^{(k+1)}$ by solving:
$\operatorname*{argmin}_{f\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\{Y_{0,i}+\gamma V_{q_{n}^{*(k)}}^{\pi}(S_{1,i})-f(q_{n}(A_{0,i},S_{0,i}))\}^{2};$
4:     update $q_{n}^{*(k+1)}:=f_{n}^{(k+1)}\circ q_{n}$;
5:     if $\|q_{n}^{*(k+1)}-q_{n}^{*(k)}\|_{P_{n}}<\varepsilon$ then
6:        set $q_{n}^{*}:=q_{n}^{*(k+1)}$;
7:        break;
8:     end if
9:  end for
10:  return  $q_{n}^{*}$;
Our novel calibration algorithm, Isotonic Bellman Calibration, is outlined in Algorithm 2, where $\mathcal{F}_{\text{iso}}$ denotes the space of all monotone non-decreasing (isotonic) functions. Isotonic Bellman calibration combines isotonic regression—a distribution-free calibration method widely used in prediction (Niculescu-Mizil and Caruana, 2005; Van Der Laan et al., 2023; van der Laan and Alaa, 2025)—with fitted $Q$-iteration for solving the Bellman equation (Munos and Szepesvári, 2008). Since isotonic regression solutions are generally non-unique, we follow Groeneboom and Lopuhaa (1993) and select the unique càdlàg, piecewise constant solution with jumps only at observed values of $q_{n}$. When paired with fitted $Q$-iteration, isotonic regression acts as a data-driven histogram estimator that bins the one-dimensional space $\{q_{n}(a,s):a\in\mathcal{A},\,s\in\mathcal{S}\}$. Upon convergence, the calibrated estimator $q_{n}^{*}(a,s)$ is evaluated as the empirical mean of the Bellman outcome $Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})$ over observations where $q_{n}(A_{0,i},S_{0,i})$ falls into the same bin as $q_{n}(a,s)$, thereby ensuring empirical calibration as defined in (10).
Since calibration involves additional fitting, it is important that the initial Q-function estimator $q_{n}$ in Algorithm 2 is obtained from a dataset independent of the data used for calibration. In practice, this can be achieved via sample splitting, wherein one half of the data is used to estimate $q_{n}$, and the other half is used to calibrate $q_{n}$ and compute the plug-in estimator of the linear functional. To improve data efficiency, cross-fitting techniques for nuisance estimation can be employed (van der Laan et al., 2011; Chernozhukov et al., 2018a). We provide a cross-fitted variant of Algorithm 2 in Appendix C.
5.3 Asymptotic theory
In this section, we show that the estimator $\psi_{n}^{*}$ is asymptotically linear and superefficient for $\Psi$, while achieving nonparametric efficiency for the oracle parameter $\Psi_{q_{0}}$.
Our main result of this section is the following theorem. We discuss its conditions in Appendix C.4. We introduce the following notation. Let $\varphi_{n,q_{n}^{*}}$ denote the map $(s,a,y,s^{\prime})\mapsto m(s,a,q_{n}^{*})+\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(a,s)\{y+\gamma q_{n}^{*}(a,s^{\prime})-q_{n}^{*}(a,s)\}-\psi_{n}^{*}$, and let $\varphi_{0,q_{0}}$ denote $(s,a,y,s^{\prime})\mapsto m(s,a,q_{0})+\mathcal{T}_{0,q_{0}}(\alpha_{0,q_{0}})(a,s)\{y+\gamma V^{\pi}(q_{0})(s^{\prime})-q_{0}(a,s)\}-\Psi_{q_{0}}(P_{0})$.
@@ 5.3 Asymptotic theory
If, in addition, D1 holds in a Hellinger neighborhood of $P_{0}$, then $\varphi_{0,q_{0}}$ is the EIF for $\Psi_{q_{0}}$, and:
@@
As a consequence of Theorem 7, the calibrated estimator $\psi_{n}^{*}$ satisfies the asymptotic expansion $\psi_{n}^{*}-\psi_{0}=(P_{n}-P_{0})\varphi_{0,q_{0}}+o_{p}(n^{-1/2})$, where $\varphi_{0,q_{0}}$ is the EIF for the oracle parameter $\Psi_{q_{0}}$. Therefore, $\sqrt{n}(\psi_{n}^{*}-\Psi(P_{0}))$ is asymptotically normal with limiting variance $\sigma_{0}^{2}:=\mathrm{Var}_{0}(\varphi_{0,q_{0}})$. Given a consistent estimator of $\sigma_{0}^{2}$, inference can be conducted using Wald-type confidence intervals and tests. Under regularity conditions, this variance can be consistently estimated by computing the empirical variance of an estimate of the influence function $\varphi_{0,q_{n}^{*}}$, which requires estimating the unknown quantity $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$. The function $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$ is known up to a one-dimensional transformation of $q_{n}^{*}$, and, by properties of isotonic regression, the calibrated estimator $q_{n}^{*}$ takes on finitely many values. Consequently, $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$ can be computed efficiently using matrix formulas for discrete Markov chains (see Appendix C.2). Alternatively, $\sigma_{0}^{2}$ can be estimated via a bootstrap procedure that resamples the calibration step while holding the initial $Q$-function estimator fixed, following van der Laan et al. (2024b), thereby avoiding additional nuisance estimation altogether.
The limiting variance of $\psi_{n}^{*}$ is equal to the efficiency bound for the oracle parameter $\Psi_{q_{0}}$ under $P_{0}$. Consequently, $\psi_{n}^{*}$ is generally superefficient for the original parameter $\Psi$, as its limiting variance may be strictly below the generalized Cramér–Rao efficiency bound for $\Psi$ under the nonparametric model. The efficiency gains of $\psi_{n}^{*}$ relative to nonparametric estimators can be substantial.In particular, the key quantity $\mathcal{T}_{0,q_{0}}(\alpha_{0,q_{0}})$ appearing in the limiting variance depends on the degree of intertemporal state overlap within the level sets of the $Q$-function, rather than across the entire state space. Specifically for the policy value estimand, the relevant term $\mathcal{T}_{0}(\alpha_{0,q_{0}})$ in the influence function $\varphi_{0,q_{0}}$ reduces to the aggregated density ratio $E_{0}[d_{0}(A_{0},S_{0})\mid q_{0}(A_{0},S_{0})]$, where $d_{0}$ denotes the state occupancy ratio defined in (2).
The superefficiency of $\psi_{n}^{*}$ comes at the cost of irregularity: it is not a regular estimator for the target parameter $\Psi$ and may exhibit non-vanishing asymptotic bias under sampling from local alternatives to $P_{0}$ in the nonparametric model (van der Laan et al., 2023, Theorem 7). Nevertheless, Theorem 7 guarantees that $\psi_{n}^{*}$ is regular and efficient for the oracle parameter $\Psi_{q_{0}}$ at $P_{0}$, and locally asymptotically equivalent to the semiparametric DRL estimator that assumes the oracle submodel $H_{q_{0}}$ a priori. Thus, even under a local alternative $P_{0,n^{-1/2}}$, the estimator $\psi_{n}^{*}$ enables valid inference for the projection $\Psi_{q_{0}}(P_{0,n^{-1/2}})$. In contrast, semiparametric DRL estimators are also irregular for $\Psi$ under the nonparametric model but lack adaptivity and may fail to achieve $\sqrt{n}$-consistency when the assumed model is misspecified. For additional discussion of this trade-off between superefficiency and irregularity, see van der Laan et al. (2023).
5.4 Related work
In the special case of a static setting ($\gamma=0$) and the average treatment effect (ATE) functional, our estimator is asymptotically equivalent to the targeted minimum loss estimator (TMLE) proposed by Benkeser et al. (2020). That approach uses an estimate of the outcome regression as a data-adaptive dimension reduction to construct a superefficient estimator. We extend this idea to the MDP setting, replacing the outcome regression with the $Q$-function. Unlike their explicit TMLE update, we achieve debiasing via isotonic calibration, avoiding additional nuisance estimation. Our method is also inspired by the calibrated DML framework of van der Laan et al. (2024c), which shows how nuisance calibration yields higher-order debiased estimators for linear functionals of the outcome regression. The objective of that paper was to construct estimators that remain asymptotically normal even when one of the nuisance functions is inconsistently estimated. While we generalize this idea to the MDP setting, our goal is distinct: we use calibration to construct a superefficient plug-in estimator without estimating the Riesz representer. In particular, whereas van der Laan et al. (2024c) calibrate both nuisances for nonparametric inference, we calibrate a single nuisance to avoid the complexity of high-dimensional min-max estimation.
@@ 6 Numerical experiments
We consider Example 2 described in Section 2.2, where the parameter $\Psi$ represents the long-term causal effect of an A/B test. Participants are randomly assigned to either a treatment group, receiving a specific intervention, or a control group, receiving an alternative intervention or no intervention. We let $A_{t}$ denote the study assignment at time $t$, where $A_{t}=Z$ almost surely, and consider the behavior policy $\pi$ that sets the treatment $A_{t}$ equal to $Z$. We write the state at time $t$ as $S_{t}=(Z,\widetilde{S}_{t})$, and write a generic realization of $S_{t}$ as $s=(z,\widetilde{s})$. In this case, the $Q$-function $q_{0}^{\pi}$ equals the V-function $(\widetilde{s},z)\mapsto\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}\mid\widetilde{S}_{0}=\widetilde{s},Z=z\right]$. Following Tran et al. (2023), our parameter of interest is the long-term ATE $E_{0}[V^{\pi}(q_{0})((1,\widetilde{S}_{0}))-V^{\pi}(q_{0})((0,\widetilde{S}_{0}))]$.
To mimic an A/B test on online platforms, we simulate a discrete-state Markov process with four state variables: engagement, churn risk, tenure, and an overlap variable parameterized by $\beta\in\mathbb{R}$. Each variable takes values in $\{0,1,2\}$. The degree of intertemporal state overlap is governed by the parameter $\beta$, with higher values of $\beta$ inducing less overlap between treatment arms and states over time. This setup allows us to assess estimator performance under varying degrees of intertemporal overlap—a key challenge in long-term causal inference. The simulation details are provided in Appendix D.2.
We estimate the $V$-function using fitted $Q$-iteration with gradient-boosted regression trees implemented in lightgbm. We compare three estimators: the Bellman-calibrated plug-in estimator from Section 5, an adaptive DRL estimator that learns a data-driven model for the $V$-function, and the nonparametric DRL estimator of Kallus and Uehara (2020) and Tran et al. (2023). The adaptive DRL approach, described in Section 3.4, constructs the model class $\mathcal{H}_{n}$ by one-hot encoding the leaf nodes of a gradient-boosted sum-of-trees model for the $V$-function (see, e.g, Section 3.1 of He et al. (2014)). The Riesz representer is estimated via min-max optimization using gradient-boosted trees; the inner maximization is approximated in closed form via ridge regression on the induced tree features. Confidence intervals for the Bellman-calibrated estimator are obtained by bootstrap, following van der Laan et al. (2024c).
@@
Figure 2 summarizes the performance of the three estimators across settings with varying intertemporal overlap ($\beta$) and discount factors ($\gamma$). Appendix D.2 provides results for the good ($\beta=0$) and moderate ($\beta=0.3$) overlap settings. Estimator performance depends critically on the degree of overlap: with good overlap, all methods exhibit low bias, small standard errors, and near-nominal coverage. As overlap deteriorates ($\beta=0.3$, $0.6$), the nonparametric estimator performs poorly—showing high variance and undercoverage—reflecting its sensitivity to inverse weighting. Subfigure (d) compares the nonparametric and Bellman-calibrated estimators under low overlap at $\gamma=0.8$. The nonparametric method exhibits greater bias, variance, and coverage error, while Bellman calibration yields substantially better coverage that approaches 95% with increasing sample size. Across all settings, Bellman calibration is the most stable, consistently achieving low bias and variance. While the tree-based adaptive DRL estimator improves over the nonparametric baseline, it remains more variable than Bellman calibration under poor overlap, likely due to the higher complexity of $\mathcal{H}_{n}$.
The discount factor $\gamma$ determines how far into the future the estimator must extrapolate to evaluate long-term causal effects. As $\gamma$ increases, estimation becomes more difficult, with greater bias and standard error observed near $\gamma=0.75$. The nonparametric estimator becomes increasingly unstable at higher $\gamma$, reflecting its sensitivity to limited intertemporal overlap. In contrast, Bellman calibration maintains low bias and variance across all values of $\gamma$, underscoring the benefits of calibration and dimension reduction in overcoming overlap limitations.
7 Conclusion
In this work, we leveraged $Q$-function calibration to develop calibrated plug-in estimators. A promising direction for future research is to extend the calibrated debiased machine learning framework of van der Laan et al. (2024c) to linear functionals of solutions to integral equations. This extension could demonstrate that calibrating both the $Q$-function and the Riesz representer yields doubly robust asymptotically linear estimators (Benkeser et al., 2017), enabling valid inference, including confidence intervals and hypothesis tests, even if either component is estimated inconsistently or at a slow rate. We leave the development of such a doubly robust inference procedure for future work.

### v4

@@ 4.3 Asymptotic theory
Suppose that C1(i) holds with $H=H_{(q_{n}^{*},q_{0})}$ (continuity of linear functional) and D1 holds. Then,
$\Psi_{q_{n}^{*}}(P_{0})-\Psi(P_{0})$ $=\langle d_{0,(q_{n}^{*},q_{0})}-d_{0,q_{n}^{*}},\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}}-q_{0})\rangle_{P_{0}}.$
If D2 also holds, then $\Psi_{q_{n}^{*}}(P_{0})-\Psi(P_{0})=O_{p}\left(\|q_{n}^{*}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}\right).$
Theorem 5 shows that the bias term $\Psi_{q_{n}^{*}}(P_{0})-\Psi(P_{0})$ vanishes provided that the optimal transformation $q_{0,q_{n}^{*}}$ of $q_{n}^{*}$ converges to $q_{0}$, and the difference $d_{0,(q_{n}^{*},q_{0})}-d_{0,q_{n}^{*}}$ converges to zero. Heuristically, since $d_{0,q_{n}^{*}}(A_{0},S_{0})=E_{0}[d_{0,(q_{n}^{*},q_{0})}(A_{0},S_{0})\mid q_{n}^{*}(A_{0},S_{0})]$, this requires that conditioning on the pair $(q_{n}^{*}(A_{0},S_{0}),q_{0}(A_{0},S_{0}))$ asymptotically conveys no more information than conditioning on either component alone. Under D2, this convergence holds whenever $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(1)$. Similar conditions for DML with estimated features are assumed in benkeser2017doubly; benkeser2020nonparametric; wang2023super; bonvini2024doubly; vanderlaan2024automatic. Condition D1 is a dimension-reduced version of C1.
We now show that the estimator $\psi_{n}^{*}$ is asymptotically linear and efficient for the oracle parameter $\Psi_{q_{0}}$, and therefore superefficient for $\Psi$. In what follows, for $q\in\{q_{n}^{*},q_{0}\}$, let $\varphi_{0,q}(s,a,y,s^{\prime}):=\mathcal{T}_{0,q}(\alpha_{0,q})(a,s)\{y+\gamma V^{\pi}(q)(s^{\prime})-q(a,s)\}+m(s,a,q)-\Psi_{q}(P_{0})$. The following conditions are analogous to those of Theorem 2 and are discussed in Appendix C.4.
@@
Moreover, $\varphi_{0,q_{0}}$ is the EIF for $\Psi_{q_{0}}$, and:
@@
By Theorem 6, the calibrated estimator $\psi_{n}^{*}$ admits the expansion $\psi_{n}^{*}-\psi_{0}=P_{n}\varphi_{0,q_{0}}+o_{p}(n^{-1/2}),$ where $\varphi_{0,q_{0}}$ is the EIF for the oracle parameter $\Psi_{q_{0}}$. Hence $\sqrt{n}(\psi_{n}^{*}-\Psi(P_{0}))$ is asymptotically normal with variance $\sigma_{0}^{2}:=P_{0}\varphi_{0,q_{0}}^{2}$, and inference may be conducted using standard Wald intervals once a consistent estimator of $\sigma_{0}^{2}$ is available. Under regularity conditions, $\sigma_{0}^{2}$ can be consistently estimated by the empirical variance of an estimated influence function $\varphi_{0,q_{n}^{*}}$, which requires estimating $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$. Since $q_{n}^{*}$ is discretely valued, this quantity can be computed efficiently using matrix formulas for finite-state Markov chains (Appendix C.2). As an alternative, $\sigma_{0}^{2}$ may be estimated via a bootstrap that resamples only the calibration step while keeping the initial $Q$-function estimator $q_{n}$ fixed; theoretical justification for this approach is given in vanderlaan2024automatic.
Since $\psi_{n}^{*}$ is efficient for the dimension-reduced parameter $\Psi_{q_{0}}$, which leverages knowledge of $q_{0}$, it is superefficient for the target parameter $\Psi$. The superefficiency of $\psi_{n}^{*}$ comes at the cost of irregularity for $\Psi$: the estimator may exhibit non-vanishing asymptotic bias under sampling from a local alternative $P_{0,hn^{-1/2}}$ to $P_{0}$, with $\sqrt{n}\{\psi_{n}^{*}-\Psi(P_{0,hn^{-1/2}})\}\underset{P_{0,hn^{-1/2}}}{\rightsquigarrow}N(0,\sigma_{0}^{2})+O(h),$ and inference is therefore not uniformly valid over the nonparametric model (van1997superefficiency; van2023adaptive). Nonetheless, accepting irregularity in exchange for superefficiency can be advantageous in practice (benkeser2020nonparametric; van2023adaptive; moosavi2023costs). Theorem 6 guarantees that $\psi_{n}^{*}$ remains regular and efficient for the oracle parameter $\Psi_{q_{0}}$ at $P_{0}$, meaning $\sqrt{n}\{\psi_{n}^{*}-\Psi_{q_{0}}(P_{0,hn^{-1/2}})\}\underset{\text{\scriptsize$P_{0,hn^{-1/2}}$}}{\rightsquigarrow}N(0,\sigma_{0}^{2})$. Thus, even under sampling from $P_{0,hn^{-1/2}}$, the estimator $\psi_{n}^{*}$ continues to yield valid asymptotic inference for the projected target $\Psi_{q_{0}}(P_{0,hn^{-1/2}})$.
@@ 5 Numerical experiments
We consider Example 2 from Section 2.2, where the parameter $\Psi$ represents the long-term causal effect of an A/B test. Following tran2023inferring, the estimand is the long-term ATE $E_{0}\!\left[V^{\pi}(q_{0})\bigl(1,\widetilde{S}_{0}\bigr)-V^{\pi}(q_{0})\bigl(0,\widetilde{S}_{0}\bigr)\right].$ To mimic an online A/B test, we simulate a discrete-state Markov process with four state variables—engagement, churn risk, tenure, and an overlap variable parameterized by $\beta$—each taking values in $\{0,1,2\}$. The parameter $\beta$ governs the degree of intertemporal overlap: larger values correspond to weaker overlap between treatment arms and downstream states. This setup enables us to assess estimator performance under varying overlap conditions, a key challenge in long-term causal inference. Full simulation details are provided in Appendix D.2.
We estimate the $V$-function using fitted $Q$-iteration with gradient-boosted regression trees implemented in lightgbm. We compare three estimators: (i) the Bellman-calibrated plug-in estimator from Section 4, (ii) an adaptive DRL estimator that learns a data-driven model class for the $V$-function, and (iii) the nonparametric DRL estimator of kallus2020double and tran2023inferring. The adaptive DRL method builds $\mathcal{H}$ by one-hot encoding the leaf nodes of a boosted tree model for the $V$-function (Section 3.1 of he2014practical). The Riesz representer is then estimated via min–max optimization, with the inner maximization approximated in closed form by ridge regression on the induced tree features. Confidence intervals for the Bellman-calibrated estimator are obtained via the bootstrap procedure of vanderlaan2024automatic.
@@
Figure 2 compares the three estimators across varying levels of intertemporal overlap ($\beta$) and discounting ($\gamma$). Results for the good ($\beta=0$) and moderate ($\beta=0.3$) overlap settings are given in Appendix D.2. Performance depends strongly on overlap: when overlap is good, all methods achieve low bias, small standard errors, and near-nominal coverage. As overlap worsens ($\beta=0.3,0.6$), the nonparametric estimator deteriorates—exhibiting high variance and undercoverage—reflecting its sensitivity to inverse weighting. Subfigure (d) highlights this at $\gamma=0.8$, where the nonparametric method shows large bias, variance, and coverage error, while Bellman calibration maintains coverage near 95% as $n$ increases. Across all settings, Bellman calibration is the most stable, consistently achieving low bias and variance. The adaptive DRL estimator improves upon the nonparametric baseline but remains more variable than Bellman calibration under poor overlap, likely due to the higher complexity of $\mathcal{H}$. Finally, as $\gamma$ increases, the nonparametric estimator becomes increasingly unstable, whereas Bellman calibration remains accurate and low-variance, demonstrating robustness in long-horizon settings.
6 Conclusion and future work
We introduced a calibrated plug-in approach for inference in dynamic settings by leveraging $Q$-function calibration. A natural direction for future research is to extend the calibrated DML framework of vanderlaan2024automatic from static to dynamic models. Such an extension could show that jointly calibrating the $Q$-function and its Riesz representer yields doubly robust inference (benkeser2017doubly), enabling valid inference even when one of the two components is inconsistently estimated or converges slowly. In particular, calibrating the $Q$-function in DRL can improve robustness to misspecification of the density ratio. For this reason, we recommend $Q$-function calibration in nonparametric DRL, even when the density ratio is used for bias correction.
Additional directions include extending the methodology to time-inhomogeneous MDPs and to correlated data settings in which multiple transitions are observed per individual. Finally, the framework extends naturally to well-posed linear inverse problems more broadly, by replacing the Bellman operator with an invertible linear operator.